Write the probability distribution of a random variable R representing numbers of red ball when 3 balls are drawn in succession without replacement from a jar containing 4 red and blue balls

1
Expert's answer
2022-03-15T05:56:04-0400

We can pull out three red balls, or two red and one blue. To find the probability of three red balls, we need to repeat and multiply three times: divide the number of red balls by the total number of balls and pull out the red ball:

"P(3\text{ red})=\frac{4}{5}\cdot\frac{3}{4}\cdot\frac{2}{3}=0.4\\
P(2\text{ red})=1-P(3\text{ red})=0.6"

We can write down the probability distribution:

"R=\begin{pmatrix}
 2 & 3 \\
 0.6 & 0.4
\end{pmatrix}"